# Sub-millisecond electric field sensing with an individual rare-earth doped ferroelectric nanocrystal

### v1

@@ 2.3 Sensitivity of the nanoBTO:Yb,Er electric field sensor
With nanoBTO:Yb,Er #3, of size 130 nm, we were able to detect a field amplitude kV/cm within s with a signal-to-noise ratio of 1.6. The minimum field detectable during an optical cycle is then kV/cm, leading to a sensitivity kV/cm/. Considering this sensitivity, if we integrate over 1 ms (a third of the total duration of a neuron action potential), we should be able to detect a field kV/cm, which is smaller than the peak-to-peak axonal membrane potential of kV/cm. Moreover, the sensitivity should improve by considering larger particles as the signal increases with more ions being exposed to the electric field.
@@ 3 Conclusions
We synthesized, by hydrothermal method, erbium and ytterbium-doped barium titanate NCs of average size 158 nm, with effective doping estimated to be 1.7% and 1.2% for Er3+ and Yb3+ respectively. Despite the doping, the nanoBTO retained its tetragonal structure to which ferro- and piezo-electric properties are associated at room temperature of 293 K. We observed that 61% of the nanocrystals displayed a variation of up-conversion intensity upon application of +10 V, -10 V or rarely for both polarities. Moreover, about half (47%) of the nanocrystals had a fast response (characteristic duration <100 s), compatible with a faithful sampling of action potential temporal profile. UC intensity changes are attributed to fast modifications of the crystal field around erbium ions, that lower the symmetry inducing larger intra-4f transitions probabilities. We suggested that the NCs which do not yield a detectable UC change, have their crystallographic axis perpendicular to the applied electric field (i.e. this axis lies in the substrate plane). Indeed, we established in a recent piezoresponse force microscopy study 35, that the PFM displacement amplitude is one order of magnitude smaller for such nanoBTO crystal axis orientation than for NCs with axis parallel to the applied electric field (i.e., perpendicular to the substrate). It is desirable to make the nanosensor more isotropic and less sensitive to the electric field orientation. As the largest crystal deformations induced by the electric field take place at the NC surface, we envision that core-shell perovskite structures (like the one reported in Ref. 37) may be grown by epitaxy in such a way that shell axis orientation is orthogonal to all facets. In such composite, the in-plane crystal deformation should be large and independent of the external field direction.
@@
With the nanoBTO:Yb,Er sensor we developed, we demonstrated a single particle sensitivity to electric field sensing of 4.8 kV/cm/. There are still some margin for improvement by enhancing erbium emission. In particular, we could incorporate lithium ions in addition to erbium and ytterbium, as it has been shown that lithium (by either A site substitution or by interstitial insertion) can increase the UC intensity by further lowering crystal symmetry around Er3+ ions 38. Furthermore, a very thin shell (thickness<5 nm) of silanized ytterbium complex could be used to enhance the absorption cross-section of the near-infrared excitation light by the nanocrystal.
@@
The application of the nanoBTO:Yb,Er to neuron cell membrane potential variation sensing will require the sensor to be brought into contact with the membrane as the Debye layer electric field shielding takes place within a few nanometers. Grafting of bioconjugates to BTO has been reported by several groups 19, 20, 21 and could be leveraged to attach molecules that are able to anchor the sensor in the cell membrane. The nanoBTO:Yb,Er electric field sensor may also find applications in other scientific areas where electric field needs to be probed with nanometer and sub-millisecond resolutions at room temperature. In these domains, the nitrogen-vacancy center in a diamond nanopillar offers a record sensitivity of only 0.24 kV/cm/, but its implementation is still complex 39.
@@ 4.2 Characterization instruments and methods
X-Ray diffraction spectrum was acquired using a Bruker D2 diffractometer using Cu K1 radiation (wavelength: 1.5418 Å) and Cu K2 radiation (wavelength: 1.5444 Å). XRD data were collected at room temperature between angles 20∘ and 60∘ by steps of 0.02∘. The EVA software was then used to determine the phase composition of the material. Transmission electron microscopy (TEM) analyses in parallel and in scanning mode (TEM - STEM) were conducted utilising a FEI Thermofisher Titan3 G2 80-300 microscope, operated at 300 kV and equipped with a Cs probe corrector, STEM detectors, an Ultrascan 1000 XP TEM camera and a SuperX detector for energy-dispersive X-ray (EDX) analyses. EDX quantifications have been performed by using the Cliff-Lorimer method 40. As Ti-Kα (4.510 keV) and Ba-Lα (4.465 keV) rays have very close energies, an optimisation of -factors has been performed using pure BTO.
@@ 4.4 Immobilization of nanoBTO
A 1% weight poly (3,4-ethylenedioxythiophene) polystyrene sulfonate (PEDOT:PSS) in water suspension was spin coated on a 170 m thick cover glass covered with a semitransparent layer of ITO (80 nm ITO layer thickness), forming a nm thick layer, on top of which the nanoBTO:Yb,Er aqueous suspension is deposited. Due to the good wettability of BTO by water 41, the concave meniscus on the edges of the NCs drags them down by surface tension, leading to a final position of the particle emerging from the polymer layer as we showed in another work 35.
@@
The authors thank Shuping Liu for the synthesis of non ferroelectric Y2O3:Er3+ nanocrystals used for the control experiment, and Céline Fiorini-Debuisschert for her help in finding a solution to firmly attach the nanocrystals to the conductive coverglass substrate. We also thank Brahim Dkhil, Aleix G. Güell, Julien Boudon, Lionel Maurizi and Nadine Millot for fruitful discussions. The AFM used in this work was purchased thanks to the FOSTER program of ENS Paris-Saclay. This work has received financial support to B.D. and F.T. from the CNRS through the MITI interdisciplinary program and from the French National Research Agency (ANR, grant numbers ANR-21-CE09-0028 and ANR-21-CE09-0033).
@@
Supporting information
Sub-millisecond electric field sensing with an
individual rare-earth doped ferroelectric
nanocrystal
A. K. Muraleedharan, J. Zou, M. Vallet, A. Zaki, C. Bogicevic, C. Paillard, K. Perronet and F. Treussart
1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | |
[…]
UC with -10 V |

### v2

@@ 2.3 Sensitivity of the nanoBTO:Yb,Er electric field sensor
With nanoBTO:Yb,Er #3, of size 130 nm, we were able to detect a field amplitude kV/cm within s with a signal-to-noise ratio of 1.6. The minimum field detectable during an optical cycle is then kV/cm, leading to a sensitivity kV/cm/. Considering this sensitivity, if we integrate over 1 ms (a third of the total duration of a neuron action potential), we should be able to detect a field kV/cm. This value is slightly smaller than the peak-to-peak axonal membrane potential variation of kV/cm during an action potential, opening prospects in the detection of such an event. Moreover, the sensitivity should improve by considering larger particles as the signal increases with more ions being exposed to the electric field.
@@ 3 Conclusions
We synthesized, by hydrothermal method, erbium and ytterbium-doped barium titanate NCs of average size 158 nm, with effective doping estimated to be 1.7% and 1.2% for Er3+ and Yb3+ respectively. Despite the doping, the nanoBTO retained its tetragonal structure to which ferro- and piezo-electric properties are associated at room temperature of 293 K. We observed that 61% of the nanocrystals displayed a variation of up-conversion intensity upon application of +10 V, -10 V or rarely for both polarities. Moreover, about half (47%) of the nanocrystals had a fast response (characteristic duration <100 s), compatible with a faithful sampling of action potential temporal profile. UC intensity changes are attributed to fast modifications of the crystal field around erbium ions, that lower the symmetry inducing larger intra-4f transitions probabilities. We suggested that the detectable UC change may be due either to some current leakage through the particle or/and to unfavorable crystallographic axis orientation relative to the applied electric field.
Mobile oxygen vacancies are considered to be responsible for current leakage in perovskite oxide ferroelectrics ABO3, strongly affecting their ferroelectric and piezoelectric properties. In rare-earth ions doped ABO3 crystals, such vacancies appear to ensure charge neutrality in circumstances when RE3+ substitute to A and/or B sites without compensating the charges. Conduction measurements on individual NC will help to state on the presence of such vacancies 32, and if it is confirmed, the synthesis will need to be further improved.
Regarding the influence of the NC orientation relative to the electric field on sensing, we established in a recent piezoresponse force microscopy study 38, that the PFM displacement amplitude is one order of magnitude smaller for NC with their axis perpendicular the applied electric field than for parallel to the field (i.e., perpendicular to the substrate). It is desirable to make the nanosensor more isotropic and less sensitive to the electric field orientation. As the largest crystal deformations induced by the electric field take place at the NC surface, we envision that core-shell perovskite structures (like the one reported in Ref. 43) may be grown by epitaxy in such a way that shell axis orientation is orthogonal to all facets. In such composite, the in-plane crystal deformation should be large and independent of the external field direction.
@@
With the nanoBTO:Yb,Er sensor we developed, we demonstrated a single particle sensitivity to electric field sensing of 4.8 kV/cm/, corresponding to the ability to distinguish a field of 152 kV/cm from the noise, with a millisecond temporal resolution. As the peak-to-peak field amplitude of 200 kV/cm (during the 1 ms duration of an action potential) remains only slightly larger than the sensor sensitivity, improving the latter is desirable. To do so we could try to enhance erbium emission. In particular, we could incorporate lithium ions in addition to erbium and ytterbium, as it has been shown that lithium (by either A site substitution or by interstitial insertion) can increase the UC intensity by further lowering crystal symmetry around Er3+ ions 44. Furthermore, a very thin shell (thickness<5 nm) of silanized ytterbium complex could be used to enhance the absorption cross-section of the near-infrared excitation light by the nanocrystal.
The application of the nanoBTO:Yb,Er to neuron cell membrane potential variation sensing requires the sensor to be brought into contact with the membrane as the Debye layer electric field shielding takes place within a few nanometers. We envision that fulfilling this localization constrain is much more critical for sensing than controlling the nanocrystal crystalline axis orientation relative to the electric field. Such membrane localization may be achieved by surface functionalization of the nanocrystals. Indeed, grafting of bioconjugates to BTO has been reported by several groups 19, 20, 21. Following similar strategies, our collaborators are currently functionalizing the nanocrystals with polyethylene glycol (PEG)-cholesterol conjugates, able to anchor the sensor in the cell membrane and at the same time, thanks to the PEG chains, prevent the adsorption of culture medium proteins that would lead to non-specific interactions.
@@
Finally, the nanoBTO:Yb,Er electric field sensor may also find applications in other scientific areas where electric field need to be probed with nanometer and sub-millisecond resolutions at room temperature. In these domains, the nitrogen-vacancy center in a diamond nanopillar offers a record sensitivity of only 0.24 kV/cm/, but its implementation is still complex 45.
@@ 4.2 Characterization instruments and methods
X-Ray diffraction spectrum was acquired using a Bruker D2 diffractometer using Cu K1 radiation (wavelength: 1.5418 Å) and Cu K2 radiation (wavelength: 1.5444 Å). XRD data were collected at room temperature between angles 20∘ and 60∘ by steps of 0.02∘. The EVA software was then used to determine the phase composition of the material. Transmission electron microscopy (TEM) analyses of Figure 1b-c was done with a JEOL JEM-2100Plus operated at 200 kV. in parallel and in scanning mode (TEM-STEM) Scanning TEM (STEM) acquisition were conducted utilising a FEI Thermofisher Titan3 G2 80-300 microscope, operated at 300 kV and equipped with a Cs probe corrector, STEM detectors, an Ultrascan 1000 XP TEM camera and a SuperX detector for energy-dispersive X-ray (EDX) analyses. EDX quantifications have been performed by using the Cliff-Lorimer method 46. As Ti-Kα (4.510 keV) and Ba-Lα (4.465 keV) rays have very close energies, an optimisation of -factors has been performed using pure BTO.
@@ 4.4 Immobilization of nanoBTO
A 1% weight poly (3,4-ethylenedioxythiophene) polystyrene sulfonate (PEDOT:PSS) in water suspension was spin coated on a 170 m thick cover glass covered with a semitransparent layer of ITO (80 nm ITO layer thickness), forming a nm thick layer, on top of which the nanoBTO:Yb,Er aqueous suspension is deposited. Due to the good wettability of BTO by water 47, the concave meniscus on the edges of the NCs drags them down by surface tension, leading to a final position of the particle emerging from the polymer layer as we showed in another work 38.
@@
The authors thank Shuping Liu and Godefroy Leménager for the synthesis of non ferroelectric Y2O3:Er and of NaYF4:Yb,Er nanocrystals, respectively, used for the control experiments, and Céline Fiorini-Debuisschert for her help in finding a solution to firmly attach the nanocrystals to the conductive coverglass substrate. We also thank Brahim Dkhil, Aleix G. Güell, Julien Boudon, Lionel Maurizi and Nadine Millot for fruitful discussions. The AFM used in this work was purchased thanks to the FOSTER program of ENS Paris-Saclay. This work has received financial support to B.D. and F.T. from the CNRS through the MITI interdisciplinary program and from the French National Research Agency (ANR, grant numbers ANR-21-CE09-0028 and ANR-21-CE09-0033).